# Process Supervision-Guided Policy Optimization for Code Generation

### v1

@@ Abstract
Reinforcement learning (RL) with unit test feedback has enhanced large language models’ (LLMs) code generation, but relies on sparse rewards provided only after complete code evaluation, limiting learning efficiency and incremental improvements. When generated code fails all unit tests, no learning signal is received, hindering progress on complex tasks. To address this, we propose a Process Reward Model (PRM) that delivers dense, line-level feedback on code correctness during generation, mimicking human code refinement and providing immediate guidance. We explore various strategies for training PRMs and integrating them into the RL framework, finding that using PRMs both as dense rewards and for value function initialization significantly boosts performance. Our approach increases our in-house LLM’s pass rate from 28.2% to 29.8% on LiveCodeBench and from 31.8% to 35.8% on our internal benchmark. Our experimental results highlight the effectiveness of PRMs in enhancing RL-driven code generation, especially for long-horizon scenarios.
11footnotetext: Equal contribution. This work was conducted during Ning Dai’s internship at ByteDance Inc.
@@ 1 Introduction
The rapid advancement of large language models (LLMs) has revolutionized code generation, enabling models to achieve near-human performance on programming tasks (Chen et al., 2021; Li et al., 2022; OpenAI, 2023). These models have demonstrated remarkable abilities to generate syntactically correct and functionally viable code snippets, significantly aiding software development processes. Building upon these successes, recent research has explored the use of reinforcement learning (RL) from unit test feedback to further enhance the code generation capabilities of LLMs (Le et al., 2022; Shojaee et al., 2023; Liu et al., 2023; Dou et al., 2024). By incorporating unit tests as a reward mechanism, these methods aim to guide LLMs toward generating code that not only compiles but also passes specified test cases, thereby improving overall code reliability and quality.
@@
In this work, we conduct a comprehensive analysis of how PRMs can be integrated into RL training for code generation. We explore various strategies for training a robust code PRM and investigate different methods of utilizing PRMs to improve code generation performance. Based on our experiments, we provide a practical recipe for successfully using PRMs and integrating them into RL training in the context of code generation problems. Notably, one of our key findings is that using PRMs concurrently as both dense rewards and value function initialization in RL training leads to a significant performance improvement. Our contributions can be summarized as follows:
  • We propose an effective approach that automatically generates process-level supervision data by identifying the first error line in generated code using binary search. We then train a PRM on this data to generate dense signals during RL training. To the best of our knowledge, we are the first to demonstrate that PRMs can benefit RL from unit test feedback in code generation.
  • We conduct systematic experiments to determine how to properly and effectively integrate PRMs into RL. Our analysis explores various strategies for training a high-quality code PRM and utilizing PRMs to improve code generation. We summarize our findings into a practical recipe for successfully using PRMs in the context of code generation.
  • By following the recipe, we significantly improve our in-house proprietary LLM’s pass rate from 28.2% to 29.8% on the LiveCodeBench dataset and from 31.8% to 35.8% on our in-house benchmark. Besides, we find that integrating PRMs into RL training benefits code generation in long-horizon scenarios.
@@ 2.1 Baseline Method: Reinforcement Learning from Unit Test Feedback
$R_{\texttt{UT}}(\mathbf{x},\mathbf{y})=\begin{cases}1,&\text{if the program $\mathbf{y}$ passes all unit test cases}\\ 0,&\text{otherwise}\end{cases}$
@@
3 Process Supervision-Guided Policy Optimization
Figure 1: Overview of our method. The approach consists of two main components: (1) A binary search-based method to automate PRM training data labeling, which is used to train a code PRM; and (2) Integration of the PRM into RL training as both dense rewards and value function initialization.
@@ 3 Process Supervision-Guided Policy Optimization
Motivated by this observation, we propose Process Supervision-Guided Policy Optimization, a method that integrates fine-grained feedback into the RL framework. Figure 1 illustrates the overview of our approach. By providing intermediate rewards that assess the correctness of partial code sequences, our approach guides the model more effectively toward generating functionally correct programs. This is achieved through a Process Reward Model (PRM) (Lightman et al., 2023) that evaluates each code generation step, offering dense reward signals that addresses the limitations of sparse end-of-trajectory rewards.
@@ 3.1.1 Data Collection
Figure 2: Binary search over code steps at line level to label prefixes. The first midpoint at $m=3$ is accepted, so the search interval moves to $[4,5]$. The next midpoint at $m=4$ is rejected, indicating unrecoverable errors occur after step 3.
Input: Prompt $\mathbf{x}$, response $\mathbf{y}=(\mathbf{y}^{(1)},\ldots,\mathbf{y}^{(T)})$, policy $p_{\theta}$, unit tests $\mathcal{U}$, number of completions $K$
Output: Labels $l_{t}$ for each prefix $\mathbf{y}^{\leq t}$
Initialize lower bound $L\leftarrow 1$, upper bound $R\leftarrow T$, failure point $F\leftarrow T+1$;
while $L\leq R$ do
       Compute midpoint $m\leftarrow\left\lfloor\dfrac{L+R}{2}\right\rfloor$;
       Set success flag $S\leftarrow\text{False}$;
       for $k=1$ to $K$ do
             Generate completion $\mathbf{c}_{k}\sim p_{\theta}(\cdot\mid\mathbf{y}^{\leq m})$;
             Form full program $\mathcal{P}_{k}\leftarrow(\mathbf{y}^{\leq m},\mathbf{c}_{k})$;
             Execute $\mathcal{P}_{k}$ with unit tests $\mathcal{U}$;
             if $\mathcal{P}_{k}$ passes all unit tests then
                  Set $S\leftarrow\text{True}$; break;
      if $S=\text{True}$ then  $L\leftarrow m+1$ ;
       else  $F\leftarrow m$, $R\leftarrow m-1$ ;
for $t=1$ to $T$ do
       if $t<F$ then $l_{t}\leftarrow+1$;
       else $l_{t}\leftarrow-1$;
Algorithm 1 Binary Search for Labeling Partial Code Prefixes
To effectively train the PRM, we require a dataset that provides fine-grained annotations indicating the correctness of partial code sequences. However, manually annotating the correctness of each line of code generated by the model is costly and not scalable. Instead, we employ an automated approach inspired by techniques used in recent works (Wang et al., 2024a; b; Luo et al., 2024). Our method leverages the model’s own capabilities to generate completions for partial code prefixes and uses automated testing to assess their correctness. The key idea is to determine whether a partial code prefix can be extended—by any means—into a complete program that passes all unit tests. If so, we consider the prefix as potentially correct; otherwise, it is labeled as incorrect.
@@ 3.1.2 Training
This regression formulation allows the PRM to estimate the likelihood that a given prefix can be successfully completed. Notably, aside from employing Mean Squared Error (MSE) loss, we also experimented with Cross-Entropy Loss and empirically found that MSE loss yielded better performance in our case.
@@ 3.2 Integrating PRM into RL Training
Given a learned PRM, we aim to identify best practices for enhancing code generation during RL training. While prior work Lightman et al. (2023); Wang et al. (2024a); Jiao et al. (2024); Wang et al. (2024b); Luo et al. (2024) uses PRMs to verify intermediate steps in mathematical tasks, this approach is less relevant for code generation. In mathematical domains, LLMs may generate correct answers with faulty reasoning Lightman et al. (2023), making intermediate verification essential. However, in code generation, problems are typically accompanied by multiple unit tests, making it improbable for incorrect code to pass all tests. As a result, the emphasis on intermediate verification is less applicable. Instead, we propose leveraging PRMs as auxiliary sources of dense signals to facilitate better exploration during RL training. While preliminary attempts have been made to incorporate PRMs into RL trainingWang et al. (2024a), these efforts are limited and warrant a more thorough investigation. We explore the following methods to integrate PRMs effectively:
@@
Similar to Wang et al. (2024a), we use PRMs to provide step-level reward signals that guide more efficient policy exploration during RL training. By rating the correctness of each line in the code response, the PRM supplies “dense” rewards that encourage the policy to explore more promising code paths, leading to improved performance.
@@
PRM as Both Dense Rewards and Value Initialization.
To fully capitalize on the advantages of PRMs, we combine both approaches. By using PRMs for dense rewards and as an initialization for the value model, we aim to enhance RL training through improved exploration and more effective credit assignment.
@@ 4.1 Experimental Setup
We utilize in-house datasets to train our model for code generation. Specifically, the training set, $\mathcal{D}_{\text{train}}$, is a comprehensive Reinforcement Learning with Human Feedback (RLHF) dataset that includes, as a subset, approximately $30,000$ diverse coding problems. Each of these problems is paired with unit tests designed to validate the functional correctness of the generated code. For evaluation, we employ two benchmarks: the publicly available LiveCodeBench( Jain et al. (2024)) and our proprietary code generation benchmark (InHouseBench). LiveCodeBench is a rigorous benchmark consisting of 612 coding tasks, collected between May 1, 2023, and September 1, 2024, designed to assess the code generation capabilities of LLMs. InHouseBench comprises 245 challenging coding problems in Chinese, spanning both Python and C++. All test problems are novel and unseen in public datasets, ensuring there is no risk of data contamination. The benchmark contains two problem categories: (1) Contest, which includes 169 coding competition-style problems, and (2) NL2Alg, which comprises 76 problems focused on translating natural language algorithm descriptions in Chinese into executable code. For evaluation, we let each model generate 10 candidate responses for each problem, using a temperature of 0.2, nucleus sampling with top-$p\text{=}0.95$, and top-k sampling with $k\text{=}128$, following common practice. We adopt Pass@1 as the evaluation metric, in line with previous work (Kulal et al., 2019; Chen et al., 2021; Jain et al., 2024).
Base Models and RL Baseline.
The base model used in our experiments is a small in-house model, referred to as Doubao-Lite throughout the remainder of this paper. Initially, Doubao-Lite was fine-tuned on our Supervised Fine-Tuning (SFT) dataset, resulting in Doubao-Lite-SFT, which then served as the initialization for the subsequent RLHF training phase. We fine-tune Doubao-Lite-SFT ($\pi_{\mathrm{ref}}$) on the RLHF dataset $\mathcal{D}_{\text{train}}$ using Proximal Policy Optimization(PPO) (Schulman et al., 2017) to obtain Doubao-Lite-RL ($\pi_{\theta}$). In our setup, two types of Outcome Reward Models (ORMs) are employed as the objective functions for RL training. For non-coding prompts, we use a general reward model, $R_{\texttt{general}}(\mathbf{x},\mathbf{y})$, derived from preference learning on a human-annotated dataset (Ouyang et al., 2022). For coding prompts, the ORM is defined as a binary indicator of whether the response passes all unit tests, $R_{\texttt{UT}}$ ($\mathbf{x}$, $\mathbf{y}$). Following  Ouyang et al. (2022), RLHF optimization objective is defined as:
@@
$\max_{\theta}\sum_{\mathbf{x}\in\mathcal{D}_{train}}\mathbb{E}_{\mathbf{y}\sim\pi_{\theta}(\mathbf{y}\mid\mathbf{x})}\left[R_{\texttt{UT}}(\mathbf{x},\mathbf{y})+\lambda R_{\texttt{PRM}}(\mathbf{x},\mathbf{y})-\beta\mathrm{KL}(\pi_{\theta}\parallel\pi_{\mathrm{ref}})\right],$ (3)
@@
4.2 Key Considerations for Integrating PRM into RL Training
@@ 4.2.1 PRM Training: More Data or Better Data?
Recent research on LLMs highlights that data quality often outweighs quantity (Gunasekar et al., 2023; Li et al., 2023b). We found the same holds true for PRM training data selection. Although automated data annotation allows for the generation of large volumes of PRM training data through model sampling, our experiments showed that increasing data volume can sometimes degrade PRM performance when integrated into RL. In contrast, a smaller, carefully curated subset of the full dataset led to better supervision and improved outcomes. For example, when all sampled responses to a given prompt either consistently pass (or fail) unit tests, PRM gains little useful information. In such cases, the model can only learn to predict the correct (or incorrect) label when it encounters the same prompt again, limiting its ability to generalize. We explored various strategies for selecting and filtering data, as detailed in Section 4.3.
@@ 4.2.2 RL Training: Alleviating PRM Hacking
As described in Section 4.1, when used to provide dense rewards, PRM assigns line-level reward signals at the end-of-line tokens in the LLM-generated response. However, if we directly use the predictions of the learned PRM, denoted as $R_{\phi}$, as the reward signal $R_{\mathrm{PRM}}$ in 3, we observed that this can be exploited. Specifically, the policy may generate numerous lines for which PRM predicts positive rewards, thus inflating the overall reward. This occurs because writing more lines allows the model to accumulate excessive intermediate rewards, effectively hacking the optimization objective. To mitigate this issue, we apply length normalization to the PRM predictions. Given a prompt $\mathbf{x}$ and a response $\mathbf{y}$ with $T$ lines, $\mathbf{y}$ = ($\mathbf{y}^{(1)}$, $\mathbf{y}^{(2)}$, …, $\mathbf{y}^{(T)}$), we define the PRM dense reward signal at the $m$-th line as:
$R_{\mathrm{PRM}}(\mathbf{y}^{(m)})=\frac{1}{T}\cdot R_{\phi}(\mathbf{x},\mathbf{y}^{\leq m}).$
This normalization ensures that the policy does not gain higher cumulative rewards by generating trivial or unnecessarily long responses, as the accumulated reward is bounded within the range of $[-1,1]$ regardless of the response length.
Assigning additional neutral label into PRM training.
While length normalization helps reduce PRM exploitation, it is insufficient to fully prevent PRM hacking. Even with normalization, we empirically observed that the model can still exploit PRM by generating excessive comment lines within the code. The underlying issue is that writing a correct comment is often far easier than producing correct code. As a result, the model can artificially inflate the PRM reward by including unnecessary comment lines. To address this issue, we introduce an additional neutral label in the PRM annotation, as defined in Equation 1:
@@
By assigning a neutral label (0) to comment lines, we remove the reward bias that encourages the model to generate unnecessary comments. This adjustment ensures that only meaningful contributions to the code, rather than extraneous comments, are rewarded by the PRM.
@@
Dense Value LiveCodeBench InHouseBench
Reward Init. Easy Medium Hard Overall Contest NL2Alg Overall
GPT-4o-mini - - 81.9 27.2 3.6 40.7 43.8 68.4 51.4
Qwen2-72B - - 65.0 21.3 2.8 32.2 14.8 51.3 26.1
Gemini-Flash-1.5 - - 67.7 13.1 1.9 29.6 - - -
DeepseekCoder-33B - - 60.8 14.8 1.2 27.7 10.3 50.3 22.7
Ours-SFT - - 55.3 9.3 0.3 23.5 10.4 41.4 20.0
Ours-RL × × 70.0 7.2 1.7 28.2 24.4 48.7 31.8
× 67.9 8.9 1.9 28.2 25.0 45.4 31.4
× 68.5 9.9 2.5 28.9 25.2 48.1 32.3
69.3 12.0 1.6 29.8 27.9 53.5 35.8
Table 1: Comparison of model performance (Pass@1) across LiveCodeBench and InHouseBench with PRM as DenseReward and ValueInit settings. The performance of our models (InHouse-Lite series) on both LiveCodeBench and InHouseBench are averaged over 10 independent runs. We also report the performance of other public models for comparative purpose. The performance of Gemini-Flash-1.5 on InHouseBench is omitted due to legal considerations.

### v2

@@ Abstract
Reinforcement learning (RL) with unit test feedback has enhanced large language models’ (LLMs) code generation, but relies on sparse rewards provided only after complete code evaluation, limiting learning efficiency and incremental improvements. When generated code fails all unit tests, no learning signal is received, hindering progress on complex tasks. To address this, we propose a Process Reward Model (PRM) that delivers dense, line-level feedback on code correctness during generation, mimicking human code refinement and providing immediate guidance. We explore various strategies for training PRMs and integrating them into the RL framework, finding that using PRMs both as dense rewards and for value function initialization significantly boosts performance. Our experimental results also highlight the effectiveness of PRMs in enhancing RL-driven code generation, especially for long-horizon scenarios.
@@ 1 Introduction
The rapid advancement of large language models (LLMs) has revolutionized code generation, enabling models to achieve near-human performance on programming tasks (Chen et al., 2021a; Li et al., 2022; OpenAI, 2023). These models have demonstrated remarkable abilities to generate syntactically correct and functionally viable code snippets, significantly aiding software development processes. Building upon these successes, recent research has explored the use of reinforcement learning (RL) from unit test feedback to further enhance the code generation capabilities of LLMs (Le et al., 2022; Shojaee et al., 2023; Liu et al., 2023; Dou et al., 2024). By incorporating unit tests as a reward mechanism, these methods aim to guide LLMs toward generating code that not only compiles but also passes specified test cases, thereby improving overall code reliability and quality.
@@
In this work, we conduct a comprehensive analysis of how PRMs can be integrated into RL training for code generation. We explore various strategies for training a robust code PRM and investigate different methods of utilizing it to improve code generation performance. Based on our experiments, we provide a practical recipe for successfully using PRMs and integrating them into RL training in the context of code generation. Notably, one of our key findings is that using PRMs concurrently as both dense rewards and value function initialization in RL training improves learning efficiency and leads to a significant performance improvement. Our contributions can be summarized as follows:
  • We present a practical and effective pipeline that automatically generates process-level supervision data, trains a PRM based on these supervisions, and integrates it into RL training to provide dense feedback signals. To the best of our knowledge, we are the first to demonstrate that PRMs can benefit RL from unit test feedback in code generation.
  • We systematically investigate how to best integrate PRMs into RL training, exploring different strategies for training high-quality code PRMs and utilizing them to enhance code generation. Our findings are distilled into a practical guideline for effectively incorporating PRMs in RL for code generation.
  • Following our proposed methodology, we demonstrate significant improvements in pass rates across HumanEval, MBPP, and LiveCodeBench benchmarks for two baseline LLMs. Additionally, we highlight key insights: (1) the synergy between dense rewards and value initialization maximizes PRM effectiveness, (2) PRMs encourage exploration and improve learning efficiency, and (3) PRMs enhance code generation, particularly in long-horizon scenarios.
@@ 2.1 Baseline Method: Reinforcement Learning from Unit Test Feedback
$R_{\texttt{UT}}(\mathbf{x},\mathbf{y})=\begin{cases}1,&\text{if the program $\mathbf{y}$ passes all unit tests}\\ 0,&\text{otherwise}\end{cases}$
@@
3 Process Supervision-Guided
 Policy Optimization
Figure 1: Overview of our method. Our approach consists of two main components: (1) a binary search-based method for automating PRM training data labeling, used to train a code PRM; and (2) the integration of PRM into RL training, where it serves as (a) the initialization for the value model and (b) an evaluator assessing the correctness of each line of code, providing dense reward signals.
@@ 3 Process Supervision-Guided
Motivated by this observation, we propose Process Supervision-Guided Policy Optimization (PSGPO), a method that integrates fine-grained feedback into the RL framework. Figure 1 illustrates the overview of our approach. By providing intermediate rewards that assess the correctness of partial code sequences, our approach guides the model more effectively toward generating functionally correct programs. This is achieved using a Process Reward Model (PRM) (Lightman et al., 2023), which serves as 1) the initialization for the value model and 2) an evaluator for each code generation step, providing dense reward signals that address the limitations of end-of-trajectory rewards.
@@ 3.1.1 Data Collection
To effectively train the PRM, we require a dataset that provides fine-grained annotations indicating the correctness of partial code sequences. However, manually annotating the correctness of each line of code generated by the model is costly and not scalable. Instead, we employ an automated approach inspired by techniques used in recent works (Wang et al., 2024a, b; Luo et al., 2024). Our method leverages the model’s own capabilities to generate completions for partial code prefixes and uses automated testing to assess their correctness. The key idea is to determine whether a partial code prefix can be extended—by any means—into a complete program that passes all unit tests. If so, we consider the prefix as potentially correct; otherwise, it is labeled as incorrect.
@@
Figure 2: Binary search over code steps at line level to label prefixes. The first midpoint at $m=3$ is accepted, so the search interval moves to $[4,5]$. The next midpoint at $m=4$ is rejected, indicating errors occur after step 3.
@@
Algorithm 1 Binary Search for Labeling Code Prefixes
0:  Prompt $\mathbf{x}$, response $\mathbf{y}=(\mathbf{y}^{(1)},\ldots,\mathbf{y}^{(T)})$, policy $p_{\theta}$, unit tests $\mathcal{U}$, number of completions $K$
0:  Labels $l_{t}$ for each prefix $\mathbf{y}^{\leq t}$
1:  Initialize lower bound $L\leftarrow 1$, upper bound $R\leftarrow T$, failure point $F\leftarrow T+1$
2:  while $L\leq R$ do
3:     Compute midpoint $m\leftarrow\left\lfloor(L+R)/2\right\rfloor$
4:     Set success flag $S\leftarrow\text{False}$
5:     for $k=1$ to $K$ do
6:        Generate completion $\mathbf{c}_{k}\sim p_{\theta}(\cdot\mid\mathbf{y}^{\leq m})$
7:        Form full program $\mathcal{P}_{k}\leftarrow(\mathbf{y}^{\leq m},\mathbf{c}_{k})$
8:        Execute $\mathcal{P}_{k}$ with unit tests $\mathcal{U}$
9:        if $\mathcal{P}_{k}$ passes all unit tests then
10:           Set $S\leftarrow\texttt{True};$break
11:        end if
12:     end for
13:     $L\leftarrow\text{if }\;S=\texttt{True}\;\text{ then }\;m+1\text{ else }L$
14:     $F\leftarrow\text{if }\;S=\texttt{True}\;\text{ then }\;F\;\text{ else }m$
15:     $R\leftarrow\text{if }\;S=\texttt{True}\;\text{ then }\;R\;\text{ else }m-1$
16:  end while
17:  for $t=1$ to $T$ do
18:     $l_{t}\leftarrow\text{if }t<F\text{ then }+1\text{ else }-1$
19:  end for
@@ 3.1.2 Training
This regression formulation allows the PRM to estimate the likelihood that a given prefix can be successfully completed. Notably, we also experimented with Cross-Entropy loss and empirically found that MSE loss yielded better performance.
@@ 3.2 Integrating PRM into RL Training
Given a learned PRM, we aim to identify best practices for enhancing code generation during RL training. While prior work has used PRMs to verify intermediate steps in mathematical tasks (Lightman et al., 2023; Wang et al., 2024a; Jiao et al., 2024; Wang et al., 2024b; Luo et al., 2024), their potential for guiding code generation remains largely unexplored. In mathematical domains, LLMs may generate correct answers with faulty reasoning (Lightman et al., 2023), making intermediate verification essential. However, in code generation, problems are typically accompanied by multiple unit tests, making it improbable for incorrect code to pass all tests. As a result, the emphasis on intermediate verification is less applicable. Instead, we propose leveraging PRMs as auxiliary sources of dense signals to facilitate better exploration during RL training. While preliminary attempts have been made to incorporate PRMs into RL training (Wang et al., 2024a), these efforts are limited and warrant a more thorough investigation. We explore the following methods to integrate PRMs effectively:
@@
Similar to Wang et al. (2024a), we use PRMs to provide step-level reward signals that guide more efficient policy exploration during RL training. By rating the correctness of each line in the code response, the PRM supplies “dense” rewards that encourage the policy to explore more promising code paths.
@@ 4.1 Experimental Setup
We utilize in-house datasets to train our model for code generation. Specifically, the training set, $\mathcal{D}_{\text{train}}$, is a comprehensive Reinforcement Learning with Human Feedback (RLHF) dataset that includes, as a subset, approximately $30,000$ diverse coding problems. Each of these problems is paired with unit tests designed to validate the functional correctness of the generated code. For evaluation, we use three widely adopted benchmarks: HumanEval (Chen et al., 2021b), which contains 164 hand-crafted programming problems; MBPP (Austin et al., 2021), a dataset of 974 crowd-sourced Python programming problems (where problems with IDs 11–510 are used for evaluation); and LiveCodeBench (Jain et al., 2024), a comprehensive benchmark designed to assess the code generation capabilities of LLMs. Among the various releases of LiveCodeBench, we use LiveCodeBench v3, which includes 612 coding tasks collected between May 2023 and July 2024. In our experiments, we restricted LLMs to a single-turn chat completion setting. For each coding problem, we directly input the problem to the model without using few-shot prompting. We generate 10 candidate responses for each problem, using a temperature of 0.2, nucleus sampling with top-$p\text{=}0.95$, and top-k sampling with $k\text{=}128$, following common practice. We adopt Pass@1 as the evaluation metric, in line with previous work (Kulal et al., 2019; Chen et al., 2021a; Jain et al., 2024).
Base Models.
In our experiments, we employ two different base models, Qwen2.5-7B and Doubao-Lite, to evaluate the effectiveness of our proposed method. Qwen2.5-7B (QwenTeam, 2024) is a publicly released causal language model from the Qwen series (Yang et al., 2024), recognized for its strong performance across diverse domains. Doubao-Lite is an in-house model of comparable size and capability to Qwen2.5-7B but utilizes a different network architecture.
SFT and RL Baseline.
Initially, both Qwen2.5-7B and Doubao-Lite are fine-tuned on our Supervised Fine-Tuning (SFT) dataset, yielding Qwen2.5-7B-SFT and Doubao-Lite-SFT, which serve as the starting points for the subsequent RLHF training phase. We then further optimize these SFT models ($\pi_{\mathrm{ref}}$) on the RLHF dataset $\mathcal{D}_{\text{train}}$ using Proximal Policy Optimization (PPO) (Schulman et al., 2017), resulting in the RL models Qwen2.5-7B-RL and Doubao-Lite-RL ($\pi_{\theta}$). In our setup, two types of Outcome Reward Models (ORMs) are employed as the objective functions for RL training. For non-coding prompts, we use a general reward model, $R_{\texttt{general}}(\mathbf{x},\mathbf{y})$, derived from preference learning on a human-annotated dataset (Ouyang et al., 2022). For coding prompts, the ORM is defined as a binary indicator of whether the response passes all unit tests, $R_{\texttt{UT}}$ ($\mathbf{x}$, $\mathbf{y}$). Following  (Ouyang et al., 2022), RLHF optimization objective is defined as:
@@
$\max_{\theta}\sum_{\mathbf{x}\in\mathcal{D}_{train}}\mathbb{E}_{\mathbf{y}\sim\pi_{\theta}(\mathbf{y}\mid\mathbf{x})}[$ $R_{\texttt{UT}}(\mathbf{x},\mathbf{y})+\lambda R_{\texttt{PRM}}(\mathbf{x},\mathbf{y})$ (3)
$\qquad-\beta\mathrm{KL}(\pi_{\theta}\parallel\pi_{\mathrm{ref}})],$
@@
4.2 Key Aspects for Integrating PRM into RL Training
@@ 4.2.1 PRM Training: More Data or Better Data?
Recent research on LLMs highlights that data quality often outweighs quantity (Gunasekar et al., 2023; Li et al., 2023b). We found the same holds true for PRM training data selection. While automated annotation enables large-scale data generation via model sampling, our experiments showed that increasing volume can sometimes degrade PRM performance in RL. Instead, a smaller, well-curated subset provided better supervision and improved outcomes. For instance, when all sampled responses to a prompt consistently pass or fail unit tests, PRM learns little beyond memorization, limiting generalization. We explored various selection and filtering strategies to mitigate this, as detailed in Section 4.3.
@@ 4.2.2 RL Training: Alleviating PRM Hacking
As described in Section 4.1, PRM provides dense rewards by assigning line-level signals at end-of-line tokens. However, directly using PRM predictions, $R_{\phi}$, as the reward signal $R_{\mathrm{PRM}}$ in 3 allows exploitation: the policy can generate excessive lines with positive PRM rewards, artificially inflating the total reward. To prevent this, we apply length normalization. Given a prompt $\mathbf{x}$ and a response $\mathbf{y}$ with $T$ lines, $\mathbf{y}$ = ($\mathbf{y}^{(1)}$, $\mathbf{y}^{(2)}$, …, $\mathbf{y}^{(T)}$), the PRM dense reward at line $m$ is:
$R_{\mathrm{PRM}}(\mathbf{y}^{(m)})=\frac{1}{T}\cdot R_{\phi}(\mathbf{x},\mathbf{y}^{\leq m}),$
which ensures rewards remain bounded in $[-1,1]$, preventing the policy from gaining an advantage by generating excessively long responses.
Neutral Labeling in PRM Training.
While length normalization curbs reward inflation, models can still exploit PRM by generating excessive comments, which are easier to write than correct code. To address this, we introduce a neutral label in PRM annotation, as an extention of Eq. (1):
@@
By assigning a neutral label (0) to comments, we eliminate the incentive to generate unnecessary comments, ensuring PRM rewards only meaningful code contributions.
Table 1: Model performance comparison (Pass@1) on HumanEval, MBPP, and LiveCodeBench datasets. The first section presents the results of public models, alongside the models used in our experiments (Qwen2.5-7B and Doubao-Lite series). The second and third sections detail the performance of RL models under different configurations of PRM usage. Notably, for both models, the best overall performance is achieved when PRM is applied for both Dense Reward and Value Initialization (ValueInit).
@@
Dense Value HumanEval MBPP LiveCodeBench
Reward Init. Easy Medium Hard Overall
GPT-4o-mini - - 87.2 71.8 81.9 27.2 3.6 40.7
Qwen2-72B - - 64.6 76.9 65.0 21.3 2.8 32.2
DeepseekCoder-33B - - 79.3 70.0 60.8 14.8 1.2 27.7
Qwen2.5-7B-SFT - - 67.8 58.1 50.7 16.5 0.9 24.9
Doubao-Lite-SFT - - 59.3 59.9 55.3 9.3 0.3 23.5
Qwen2.5-7B-RL × × 73.8 62.4 60.9 13.7 1.4 27.5
× 75.4 63.1 62.8 17.1 1.7 29.6
× 76.0 63.4 63.1 14.5 1.1 28.5
74.3 65.4 66.3 15.3 1.7 30.1
Doubao-Lite-RL × × 65.1 61.9 70.0 7.2 1.7 28.2
× 69.8 63.3 67.9 8.9 1.9 28.2
× 70.0 62.1 68.5 9.9 2.5 28.9
70.9 63.8 69.3 12.0 1.6 29.8